OpenStudy (anonymous):

8x+3y=4 -7x+5y=-34 elimination using multiplacation.

OpenStudy (anonymous):

hmm..

OpenStudy (texaschic101):

8x + 3y = 4 --->(5)8x + 3y = 4 -7x + 5y = - 34 ->(-3)-7x + 5y = - 34 ---------------- 40x + 15y = 20 (result of multiplying by 5) 21x - 15y = 102 (result of multiplying by -3) ---------------add 61x = 122 x = 122/61 x = 2 now sub 2 in for x in either of the original equations.. 8x + 3y = 4 8(2) + 3y = 4 16 + 3y = 4 3y = 4 - 16 3y = -12 y = -12/3 y = -4 check...you can use either of the original equations... -7x + 5y = -34 -7(2) + 5(-4) = -34 -14 - 20 = -34 -34 = -34 (correct) ANSWER : x = 2 and y = -4

OpenStudy (anonymous):

my gosh. thank you.

OpenStudy (texaschic101):

your welcome :)